# I/O in Machine Learning Applications on HPC Systems: A 360-degree Survey

### v1

@@ 2.2. Common ML Phases
The dataset preparation phase is critical in transforming the data into a form that is suitable for model training and inference. It is a critical step in the ML lifecycle and is commonly performed before both the training and inference phases. There are several major data transformation techniques commonly used during this phase to aid in model training and inference. These are listed below:
@@
Data Aggregation is used to collect required features from the dataset (Jesus et al., 2015).
@@
Data Normalization allows to scale the size of the data down without losing information (Ali et al., 2014).
@@
Augmenting the dataset allows for training models on a larger dataset, which can lead to better performance. The augmentation methods employed often depend on the dataset modality. For text datasets, examples of augmentations include random swap, random insertion, random deletion, and random synonym replacement (Shorten et al., 2021). In contrast, for image datasets, augmentations include flipping, cropping, and rotation (Shorten and Khoshgoftaar, 2019). The data transformation techniques used affect both data preprocessing speeds and the amount of data needing to move from the preprocessing phase to the training phase. The variety of data transformation techniques that may need to be applied to the dataset causes the choice of file format to be significant. It has a large impact on data transformation speeds as it affects the time needed to apply a given algorithm to the dataset. For instance, when performing data aggregation where the dataset is filtered based on a feature value, processing data in a columnar format can be more efficient than in a row-based format due to the data locality (Abadi et al., 2008).
@@
The training phase that occurs during the ML lifecycle can further be categorized, as seen in Fig. 3, into a training phase and an evaluation phase. The figure shows the training loop and the inference phase. The training loop is often run iteratively for a certain number of epochs until the desired accuracy is achieved. An epoch is defined as one complete pass through the entire dataset (Brownlee, 2018). It is the responsibility of the model designers to choose the number of epochs that strikes a balance between enabling the model to find meaningful relationships between the input and target datasets while avoiding excessive epochs that lead to over-fitting. Over-fitting occurs when the model becomes excessively entwined with the training data, prioritizing memorization over the ability to generalize and make accurate predictions on new, unseen data. Furthermore, because the entire dataset is read multiple times, model designers need to select the proper data format for their dataset that allows efficient data retrieval.
@@
There are three datasets known as the training, validation, and testing datasets (Kahloot and Ekler, 2021) used during the training and evaluation phases. Although each dataset possesses distinct content, they commonly use the same modality and file formats. All three datasets often require preparation (i.e., cleaning, augmentation, etc.) before they can be used in later stages of the ML lifecycle. Dataset splitting involves dividing the dataset into training, validation, and test sets. Common techniques for this partitioning include random sampling (Meng, 2013), k-fold cross-validation (Anguita et al., 2012), and stratified sampling (Liberty et al., 2016). Due to the large amounts of potential I/O and computation required to prepare the data effectively, parallelization techniques are needed to ensure proper system utilization. ML applications typically use the parallelization tools provided by their primary ML framework (e.g., PyTorch (Pytorch, 2024) and TensorFlow (TensorFlow, 2024)) to parallelize data processing. Once the data has been prepared, it can be fed to the model for training, validation, evaluation, or inference.
@@
During the training phase, data from the training dataset is read into memory and fed to the model. The order in which the data is fed to the model is largely impacted by the optimization algorithm of choice. Gradient descent algorithms are a popular set of optimization algorithms used to train ML models. Due to stochastic gradient descent (SGD) being the most popular gradient descent algorithm for large datasets (Newton et al., 2018), the algorithms we will present are batch gradient descent (BGD) (Nakama, 2009), SGD (Bottou, 2012) and mini-batch gradient descent (MBGD) (Bottou et al., 2018). SGD is a specific case of MBGD, and in the future, we will refer to both as SGD unless a distinction is necessary. Deciding which algorithm to use depends on the dataset size, the computational resources, and the underlying model’s properties. It is important to note that gradient descent algorithms vary in their mathematical properties (Smith et al., 2020). At every epoch, each of the listed algorithms feeds the entirety of the dataset to the model. However, the frequency at which the model is updated, and the order in which the dataset is fed to the model differ.
@@
BGD is the most straightforward gradient descent algorithm. This technique is computationally expensive and is often used only when the entire dataset can fit into memory. The model is trained against the whole dataset before it can perform one update (Ruder, 2017). If the model cannot cache the entire dataset in a fast storage tier, it necessitates utilizing underlying storage tiers, leading to increased I/O operations. MBGD is an algorithm where random subsets or batches are chosen and fed to the model. The batch size varies depending on the specific model being used. However, if the batch size is small, the model becomes susceptible to noise updates (updates that introduce randomness or variability in the parameter updates), while large batch sizes, in contrast, may lead to over-fitting. The SGD algorithm involves updating the model after feeding a single data point. SGD is a specific case of MBGD where the batch size is . The random nature of the data selection when using SGD makes it difficult to cache the same mini-batches or data points from previous epochs for future ones. Each epoch involves a new random sampling of the data, introducing variability in the training process. This commonly leads to an I/O pattern of large amounts of small I/O performed on many random files. When training models on an HPC platform, datasets are located on a shared Parallel File System (PFS), which is generally designed for reading large batched files rather than a large number of small files. This poses a significant challenge for PFSs (Chowdhury et al., 2019). During model training, after a certain number of epochs, it is common practice to assess the model’s performance while tuning hyperparameters using a separate validation dataset (Ripley, 1996).
@@
Finally, an evaluation phase, which uses the test dataset, is undergone to determine the model’s accuracy and effectiveness on unseen data. As the training and evaluation phases are time-intensive stages within the ML lifecycle, distribution strategies have been developed to parallelize these tasks. The distribution strategies optimize the utilization of computational resources, thereby expediting the training process and enhancing efficiency in model development. Choosing the appropriate training and evaluation distribution strategy depends on factors such as the model size, dataset size, and available computational resources. This decision is pivotal in guaranteeing optimal training performance.
2.3. Distributing the Training and Evaluation Phases
@@
Multiple strategies exist to distribute the training and evaluation phases to take advantage of the parallelism granted by HPC systems. Two common techniques used are model parallelism and data parallelism, shown in Fig. 4 and Fig. 5, respectively. When using model parallelism, a single model is distributed across workers, and each worker trains their partition of the model (Choi et al., 2023a). It is often used when the model is too large to fit into the memory of a single node. The strategy has two pair synchronization points: one before the evaluation and another after adjusting each process’s model partition. Synchronization occurs among pairs of workers and is completed across multiple steps. This allows pairs of workers who have been updated to continue training without having to synchronize with all other workers. When using data parallelism, workers are employed, and the model is replicated times across each worker which each train against a different dataset partition. After each iteration, all workers must be synchronized, and the model weights are adjusted according to the results of the backpropagation. In most cases, model parallelism’s synchronization and communication overhead far exceeds data parallelism’s (Abdulaziz Alnowibet et al., 2022). However, both model parallelism and data parallelism involve inter-device (i.e., TPU, GPU, or CPU) communication to exchange information about model parameters and gradients. This communication overhead can impact training performance. Strategies such as gradient accumulation, where model updates are aggregated over several batches before communication or gradient compression (Chen et al., 2018) can help reduce the overhead of inter-device communication. More recent training distribution strategies include hybrid parallelism (Akiba et al., 2017), and pipeline parallelism (Devlin et al., 2019b). Hybrid parallelism combines data and model parallelism, where the data and the model are partitioned across multiple devices. In contrast, pipeline parallelism places different layers of the model on separate devices, allowing for efficient overlapping of computation and communication.
@@ 2.4. Model Checkpoints
While large amounts of small reads dominate the I/O operations performed when training ML models, a consistent pattern of writes also occurs (Párraga et al., 2021). This is due to the need to checkpoint the model as it is being trained. This can significantly impact the training performance, especially if the checkpoints are done frequently. Checkpoints are important as they allow long-running training jobs to pick up where they left off in the event of node failures. In addition, it allows for the analysis of intermediate model features in relation to the training data. Many training jobs begin from an existing model checkpoint from which fine-tuning for their specific task can be accomplished. Starting from a pretrained model can reduce the time and data needed for training. When checkpointing a model, the write size depends largely on the number of parameters the ML model requires (Dey et al., 2020). Because of the increasing complexity of ML models (Villalobos et al., 2022), checkpoint sizes can be expected to grow proportionally. Most ML frameworks use bit precision to store model parameters (Sze et al., 2017), the default parameter size on both PyTorch and TensorFlow. Large language models (LLM)s such as GPT-3 and BERT large have billion (Brown et al., 2020) and million (Devlin et al., 2019a) parameters respectively. The time needed to checkpoint the model depends not only on the number of parameters but also on the file format used (Rojas et al., 2021), making it an important consideration for training performance. In addition, when using frameworks such as TensorFlow, checkpointing is done independently by each worker, or a single worker is assigned the job. This can lead to the stragglers problem due to a single process being responsible for the extra work. When using a single worker to checkpoint the model, the time required does not shrink as more nodes are added to the training. PyTorch only recently (October , in version ) added support for distributed checkpointing, allowing for the saving and loading of models to be accomplished in parallel by multiple ranks. Efforts to optimize model checkpointing include model compression (Chen et al., 2020) and asynchronous checkpointing.
@@ 3. I/O Benchmarks and Profiling
There have been many ML benchmarks proposed, such as Fathom (Adolf et al., 2016), BenchNN (Chen et al., 2012), and MLPerf (Reddi et al., 2020). However, these tools focus on identifying the computational requirements of ML workloads and not the I/O requirements (Devarajan et al., 2021). Additionally, they target cloud platforms instead of HPC systems. This leaves a need for a proper benchmark focusing on I/O that simulates ML workloads. One attempt to address this gap is DLIO, a benchmark designed to simulate I/O access patterns commonly found in Deep Learning (DL) workloads. DLIO, which has been released as MLPerf Storage Benchmarks, allows for extensive configuration, including the selection of interfaces (HDF5, TFRecord, CSV, NPZ), file access patterns (one file per process versus shared file per process), data access patterns, I/O types, and transfer buffer sizes. Scientists and researchers can take advantage of DLIO by configuring the benchmark to mimic the expected I/O patterns of their DL workloads and using the results to configure their applications effectively. DLIO also supports DALI (NVIDIA Corporation, 2024), which is capable of loading batches to the GPU through the DALI library. Current limitations include the inability to store multiple samples per file in certain file formats, the inability to store data/label pairs in separate files, and the assumption that the samples will always be 2D images. The inability to store multiple samples per file poses a challenge when comparing workloads across ML frameworks. For example, the BERT workload uses the TFRecord format, yet there is currently no PyTorch data loader capable of loading TFRecords. Transitioning to an alternative file format requires configuration modifications, resulting in a workload that diverges from the original and may not faithfully represent it. Thus, while DLIO effectively captures a broad spectrum of DL applications, comparing different frameworks on the same workload while maintaining accuracy to the original proves challenging.
@@
Cheng et al. (Cheng and Gunawi, 2021) benchmark two object storage systems: MinIO and Ceph. In addition, they benchmark three key-value databases: MongoDB, Redis, and Cassandra. Their study attempts to find optimal configurations for parameters such as storage location, storage disaggregation granularity, access pattern, and data format. Notable takeaways include the object storage systems tested were more sensitive to storage location than key-value databases. In addition, the object storage systems tested were largely impacted by the storage disaggregation granularity because only one object could be queried at a time.
@@
Many dataset modalities, such as audio, video, and images, have a surplus of public datasets to choose from. In contrast, high-quality, large-scale graph datasets are hard to find; in addition, most commonly used graph datasets are too small to emulate the datasets found in the real world. Hu et al. (Hu et al., 2021) have created Open Graph Benchmark (OGB), which aims to enhance graph ML research by providing robust and scalable graph datasets. It is capable of producing graph datasets of three different size categories: small, medium, and large, which have nodes, nodes, and nodes, respectively. OGB enables researchers to analyze large-scale graph datasets, facilitating rapid growth in distributed graph machine learning (GML).
@@
While benchmarking tools allow ML developers to emulate various ML workload I/O patterns, profiling tools which enable developers to analyze the HPC I/O stack in its entirety are crucial for pinpointing potential bottlenecks. TensorFlow offers the TensorFlowProfiler, released in version , capable of profiling host-side and GPU tracing. However, it only provides information at the TensorFlow level, not capturing lower-level I/O operations such as POSIX or STDIO. Chien et al. (Chien et al., 2020) created tf-Darshan, a fine-grained I/O profiling tool. It relies on Darshan (Snyder et al., 2016a), a highly popular HPC I/O profiler and tracer. Tf-darshan augments Darshan, enabling the runtime extraction and analysis of Darshan’s internal data structures. The runtime extraction is accomplished through an augmentation of the Darshan library. It adds extraction functions that return its internal buffers that store information, such as record counters for I/O operations to individual files. This gives ML applications the ability to perform runtime analysis, enabling I/O optimization techniques such as specialized online I/O schedulers to perform runtime optimizations. Tf-darshan also offers a visualization tool that allows users to understand the I/O performance of their ML workloads intuitively. Current limitations include exclusive TensorFlow integration and a moderate overhead of - .

### v2

@@ 3.2. Common ML Phases
The dataset preparation phase is critical in transforming the data into a form that is suitable for model training and inference. It is a critical step in the ML lifecycle and is commonly performed before both the training and inference phases. Offline dataset preparation is preprocessing that can be done before the model training or inference phase begins. In contrast, online dataset preparation is preprocessing that occurs in real-time during the inference or training phase. There are several major data transformation techniques commonly used during both offline and online dataset preparation to aid in model training and inference. These are listed below:
@@
Data Aggregation is used to summarize or reduce features from the dataset to produce meaningful statistics such as mean, median, or sum among many others (Jesus et al., 2015).
@@
Data Normalization is used to scale the data changing absolute values while preserving the relative proportions or ratios between data points (et al, 2014).
@@
Augmenting the dataset allows for training models on a larger dataset, which can lead to better performance. The augmentation methods employed often depend on the dataset modality. For text datasets, examples of augmentations include random swap, random insertion, random deletion, and random synonym replacement (Shorten et al., 2021). In contrast, for image datasets, augmentations include flipping, cropping, and rotation (Shorten and Khoshgoftaar, 2019). The data transformation techniques used affect both data preprocessing speeds and the amount of data needing to move from the preprocessing phase to the training phase. The variety of data transformation techniques that may need to be applied to the dataset causes the file format to be significant. It has a large impact on data transformation speeds as it affects the time needed to apply a given algorithm to the dataset, e.g., when performing data aggregation where the dataset is filtered based on a feature value, processing data in a columnar format (instead of row-based) can be more efficient due to the data locality (et al, 2008).
@@
The training phase that occurs during the ML lifecycle can further be categorized, as seen in Fig. 4, into a training phase and a validation phase. The figure shows the training loop and the inference phase. The training loop is often run iteratively for a certain number of epochs until the desired accuracy is achieved. An epoch is defined as one complete pass through the entire dataset (Brownlee, 2018). It is the responsibility of the model designers to choose the number of epochs that strikes a balance between enabling the model to find meaningful relationships between the input and target datasets while avoiding excessive epochs that lead to over-fitting. Over-fitting occurs when the model becomes excessively entwined with the training data, prioritizing memorization over the ability to generalize and make accurate predictions on new, unseen data. Furthermore, because the entire dataset is read multiple times, model designers need to select the proper data format for their dataset that allows efficient data retrieval.
@@
There are three datasets known as the training, validation, and testing datasets (Kahloot and Ekler, 2021) used during the training, validation, testing phases. Although each dataset possesses distinct content, they commonly use the same modality and file formats. All three datasets often require preparation (i.e., cleaning, augmentation, etc.) before they can be used in later stages of the ML lifecycle. Dataset splitting involves dividing the dataset into training, validation, and test sets. Common techniques for this partitioning include random sampling (Meng, 2013), k-fold cross-validation (Anguita et al., 2012), and stratified sampling (et al, 2016a). Due to the large amounts of potential I/O and computation required to prepare the data effectively, parallelization techniques are needed to ensure proper system utilization. ML applications typically use the parallelization tools provided by their primary ML framework (e.g., PyTorch (Pytorch, 2024) and TensorFlow (TensorFlow, 2024)) to parallelize data processing. Once the data has been prepared, it can be fed to the model for training, validation, testing or inference.
@@
During the training phase, data from the training dataset is read into memory and fed to the model. The order in which the data is fed to the model is largely impacted by the optimization algorithm of choice. Gradient descent algorithms are a popular set of optimization algorithms used to train ML models. Due to stochastic gradient descent (SGD) being the most popular gradient descent algorithm for large datasets (Newton et al., 2018), the algorithms we will present are batch gradient descent (BGD) (Nakama, 2009), SGD (Bottou, 2012) and mini-batch gradient descent (MBGD) (Bottou et al., 2018). SGD is a specific case of MBGD, and in the future, we will refer to both as SGD unless a distinction is necessary. Deciding which algorithm to use depends on the dataset size, the computational resources, and the underlying model’s properties. It is important to note that gradient descent algorithms vary in their mathematical properties (et al, 2020h). At every epoch, each of the listed algorithms feeds the entirety of the dataset to the model. However, the frequency at which the model is updated, and the order in which the dataset is fed to the model differ.
@@
BGD is the most straightforward gradient descent algorithm. This technique is computationally expensive and is often used only when the entire dataset can fit into memory. The model is trained against the whole dataset before it can perform one update (Ruder, 2017). If the model cannot cache the entire dataset in a fast storage tier, it necessitates utilizing underlying storage tiers, leading to increased I/O operations. MBGD is an algorithm where random subsets or batches are chosen and fed to the model. The batch size varies depending on the specific model being used. However, if the batch size is small, the model becomes susceptible to noise updates (updates that introduce randomness or variability in the parameter updates), while large batch sizes, in contrast, may lead to over-fitting. The SGD algorithm involves updating the model after feeding a single data point. SGD is a specific case of MBGD where the batch size is . The random nature of the data selection when using SGD makes it difficult to cache the same mini-batches or data points from previous epochs for future ones. Each epoch involves a new random sampling of the data, introducing variability in the training process. This commonly leads to an I/O pattern of large amounts of small I/O performed on many random files. When training models on an HPC platform, datasets are located on a shared Parallel File System (PFS), designed for reading large batched files rather than a large number of small files. This poses a significant challenge for PFSs (et al, 2019a).
@@
Finally, after a certain number of epochs, it is common practice to assess the model’s performance while tuning hyperparameters using a separate validation dataset (Ripley, 1996). The same validation dataset to often used after each epoch to evaluate the model. As the training and validation phases are time-intensive stages within the ML lifecycle, distribution strategies have been developed to parallelize these tasks. The distribution strategies optimize the utilization of computational resources, thereby expediting the training process and enhancing efficiency in model development. Choosing the appropriate training and validation distribution strategy depends on factors such as model size, data set size, and available computational resources. This decision is pivotal in guaranteeing optimal training performance. It should be noted that after the model has converged, there is sometimes an additional phase called the testing phase which evaluates the performance of the model on unseen data using the testing dataset.
3.3. Distributing the Training and Validation Phases
@@
Multiple strategies exist to distribute the training and validation phases to take advantage of the parallelism granted by HPC systems. Two common techniques used are model parallelism and data parallelism, shown in Fig. 5 and Fig. 6, respectively. When using model parallelism, a single model is distributed across workers, and each worker trains their partition of the model (Choi et al., 2023). It is often used when the model is too large to fit into the memory of a single node. The strategy has two pair synchronization points: one before the validation and another after adjusting each process’s model partition. Synchronization occurs among pairs of workers and is completed across multiple steps. This allows pairs of workers who have been updated to continue training without having to synchronize with all other workers. When using data parallelism, workers are employed, and the model is replicated times across each worker which each train against a different dataset partition. After each iteration, all workers must be synchronized, and the model weights are adjusted according to the results of the backpropagation. This is the parallelization strategy used in Section 5.2 to analyze the I/O patterns observed during training. In most cases, model parallelism’s synchronization and communication overhead far exceeds data parallelism’s (Abdulaziz Alnowibet et al., 2022). However, both model parallelism and data parallelism involve inter-device (i.e., TPU, GPU, or CPU) communication to exchange information about model parameters and gradients. This communication overhead can impact training performance. Strategies such as gradient accumulation, where model updates are aggregated over several batches before communication or gradient compression (et al, 2018a) can help reduce the overhead of inter-device communication.
More recent training distribution strategies include hybrid parallelism (Akiba et al., 2017), and pipeline parallelism (Devlin et al., 2019b). Hybrid parallelism combines data and model parallelism, where the data and the model are partitioned across multiple devices. In contrast, pipeline parallelism places different layers of the model on separate devices, allowing for efficient overlapping of computation and communication. Federated learning, introduced by Google in 2016, is a training distribution strategy where the model is partitioned into multiple devices which train their model partition against private data (Mammen, 2021; McMahan et al., 2023). After some specified amount of training, devices send their model gradients to a central server, which aggregates the model gradients and redistributes the results back to the devices. This is beneficial for use cases in which the data being used to train the model is sensitive, such as in healthcare or finance. The training data is no longer aggregated into one central location resulting in an increase in privacy and security. The sending and receiving of the model gradients with a central server can result in significant network communication overhead. Various optimizations can reduce communication overhead, such as intelligently selecting the nodes or devices that participate in the training, semi-asynchronous training methods, and clustered federated learning: a strategy in which groups of nodes are prioritized over others during training based on various factors such as latency or physical distance (Xu et al., 2023).
@@ 3.4. Model Checkpoints
While large amounts of small reads dominate the I/O operations performed when training ML models, a consistent pattern of writes also occurs (et al, 2021e). This is due to the need to checkpoint the model as it is being trained. This can significantly impact the training performance, especially if the checkpoints are done frequently. Checkpoints are important as they allow long-running training jobs to pick up where they left off in the event of node failures. In addition, it allows for the analysis of intermediate model features in relation to the training data. Many training jobs begin from an existing model checkpoint from which fine-tuning for their specific task can be accomplished. Starting from a pretrained model can reduce the time and data needed for training. When checkpointing a model, the write size depends largely on the number of parameters the ML model requires (et al, 2020b). Because of the increasing complexity of ML models (et al, 2022d), checkpoint sizes can be expected to grow proportionally. Most ML frameworks use bit precision to store model parameters (Sze et al., 2017), the default parameter size on both PyTorch and TensorFlow. Large language models (LLM)s such as GPT-3 and BERT large have billion (et al, 2020i) and million (Devlin et al., 2019a) parameters respectively. The time needed to checkpoint the model depends not only on the number of parameters but also on the file format used (Rojas et al., 2021), making it an important consideration for training performance. In addition, when using frameworks such as TensorFlow, checkpointing is done independently by each worker, or a single worker is assigned the job. This can lead to the stragglers problem due to a single process being responsible for the extra work. When using a single worker to checkpoint the model, the time required does not shrink as more nodes are added to the training. PyTorch only recently (October , in version ) added support for distributed checkpointing, allowing for the saving and loading of models to be accomplished in parallel by multiple ranks. Efforts to optimize model checkpointing include model compression (Chen et al., 2020) and asynchronous checkpointing. A description of the popular ML frameworks and their corresponding checkpoint APIs can be found in Section 6.1.
3.5. BERT Real-world End-to-end ML Workflow
In this section, we present the well-known real-world end-to-end ML workflow of BERT. We discuss BERT in place of GPT because GPT is a closed source, meaning the details about the exact I/O found throughout its lifecycle are not fully disclosed. The authors of BERT mention that it has the most comparable existing pre-training method to OpenAI GPT. The information presented in this section is based on both the original BERT paper and the updated open-source GitHub repository, which specifies important data management attributes such as file formats and pre-training times. A more detailed analysis of the BERT training I/O patterns can be found in Section 5.2 where we simulate the BERT workload using the DLIO (et al, 2021c) benchmark.
Data Generation
The data used to pre-train the model is the BooksCorpus dataset (Zhu et al., 2015), which consists of books with a total of million words, and the English Wikipedia consisting of million words.
Offline Data Preparation
BERT is an unsupervised LLM, meaning it is trained on a large quantity of publicly available plain text data. The Wikipedia data is stripped of all texts other than text passages, ignoring lists, tables, and headers. The text from both datasets is processed into a set of input token sequences, which can then be fed to the model during pre-training. The input datasets are transformed into a single large plain text file, referred to as the corpus, with each sentence being separated by a new line. It is noted that the newline-delimited text sequences contained within the corpus should be actual sentences to enable “next sentence prediction.” Token segmentation is then used to divide the text into individual tokens using a subword tokenizer named WordPiece (Song et al., 2021). A distinguishing factor of BERT is that it uses bidirectional contextual representations, i.e., a given word’s representation is based on the words found to its left and right. This is accomplished by masking out a random % of the tokens within each token sequence. The masked-out tokens are then used as the prediction target for the sequence. The token sequence representation used is of the form <Question, Answer>, allowing for both a single sentence and a pair of sentences. This is done by placing special tokens such as [SEP] representing a sentence delimiter and [CLS] representing a token sequence starting point. The [SEP] token delimiter is used to separate different segments of text (e.g., question and answer pairs). The output of the data generation is a set of TFRecord files containing the token sequences.
Pre-training
The Tensorflow framework is used to train the model during pre-training. This step in the workflow takes around - days when using to Cloud TPUs. However, it is a one-step procedure from which specific versions of the models can be based on during fine-tuning. Notable hyperparameters used to pre-train BERT in the paper are a batch size of sequences where each sequence is tokens, resulting in tokens per batch. During pre-training, random sentences are chosen in pairs A and B in which % of the time sentence B is the sentence following A. This is done to enable “next sentence prediction.”
Fine-tuning
The fine-tuning results presented are relatively inexpensive and can be replicated in at most one hour on a single TPU or a few hours in a single GPU. All results presented in the paper were fine-tuned on a single TPU due to memory constraints. To begin, the model is loaded from a previously generated checkpoint generated from the pre-training phase. The checkpoint consists of three files containing the model weights. The hyperparameters used for fine-tuning are similar to pre-training other than batch size, learning rate, and number of epochs. The paper recommends training for , , or epochs using a batch size of or and a learning rate of d-, e-, or e-. Training with long sequence lengths in constrained memory environments can result in out-of-memory errors because the attention method’s memory requirement increases by the square of the sequence length.
Evaluation
After a series of pre-training and training steps, an optional validation or testing phase can be enabled or disabled through the -do_eval parameter. This phase is used to track model performance on validation or testing datasets, fine-tune hyperparameters, or select the best model to checkpoint.
Inference
When using BERT for inference, data preprocessing steps similar to those described above must be done. During either offline or online inference, text input must be tokenized and processed into token sequences. The distinguishing factor of online inference is that the token sequences are fed to the model, and their associated predictions must be processed in real-time. Due to the wide range of possible deployment scenarios, categorizing the data management and I/O that occur in this phase is challenging. However, notable factors distinguishing this phase are the need to store or stream the prediction generated from the model and that there is no longer a need to checkpoint the model.
@@ 4. I/O Benchmarks and Profiling
Many ML benchmarks have been proposed, such as Fathom (et al, 2016b) and BenchNN (et al, 2012). However, these tools focus on identifying the computational requirements of ML workloads and not the I/O requirements (et al, 2021c). In addition, they target cloud platforms. This leaves a need for a proper benchmark focusing on I/O that simulates ML workloads.
@@
One attempt to address this gap is DLIO (et al, 2021c), a benchmark designed to simulate I/O access patterns commonly found in Deep Learning (DL) workloads. DLIO, which has been released as MLPerf Storage Benchmarks (et al, 2020j), allows for extensive configuration, including the selection of interfaces (HDF5, TFRecord, CSV, NPZ), file access patterns (one file per process versus shared file per process), data access patterns, I/O types, and transfer buffer sizes. Scientists and researchers can take advantage of DLIO by configuring the benchmark to mimic the expected I/O patterns of their DL workloads and using the results to configure their applications effectively. DLIO also supports DALI (NVIDIA Corporation, 2024), which is capable of sending batches to the GPU for preprocessing and model training, primarily for image or video datasets. Current DLIO limitations include the inability to store multiple samples per file in certain file formats, the inability to store data/label pairs in separate files, and the assumption that the samples will always be 2D images. The inability to store multiple samples per file poses a challenge when comparing workloads across ML frameworks. For example, the BERT workload uses the TFRecord format, but currently, there is no PyTorch data loader capable of loading TFRecords. Transitioning to an alternative file format requires configuration modifications, resulting in a workload that diverges from the original and may not faithfully represent it. Thus, while DLIO effectively captures a broad spectrum of DL applications, comparing different frameworks on the same workload while maintaining accuracy proves challenging.
@@
Cheng et al. (Cheng and Gunawi, 2021) benchmark two object storage systems (MinIO and Ceph) and three key-value databases (MongoDB, Redis, and Cassandra). Their study attempts to find optimal configurations for parameters such as storage location, storage disaggregation granularity, access pattern, and data format. Notable takeaways are that the object storage systems tested were more sensitive to storage location than key-value databases. In addition, the object storage systems tested were impacted mainly by the storage disaggregation granularity as only one object could be queried at a time.
@@
Many dataset modalities, such as audio, video, and images, have a surplus of public datasets to choose from. In contrast, high-quality, large-scale graph datasets are hard to find; in addition, most commonly used graph datasets are too small to emulate the datasets found in the real world. Hu et al. (et al, 2021i) have created Open Graph Benchmark (OGB), which aims to enhance graph ML research by providing robust and scalable graph datasets. It is capable of producing graph datasets of three different size categories: small, medium, and large, which have nodes, nodes, and nodes, respectively. OGB enables researchers to analyze large-scale graph datasets, facilitating rapid growth in distributed graph machine learning (GML).
iBench (Brewer et al., 2020b) is a possible solution for distributed inference simulation and benchmarking and aims to provide several key metrics, including but limited to throughput, latency, ingest rate/bandwidth, pre-processing time, and GPU efficiency. Compared to MLPerf, they note two primary advantages: (1) the ability to measure distributed inference performance and (2) a more realistic performance measure. The simulation places a stream load balancer in front of the HPC inference servers. Client requests are then simulated and scheduled to run via the load balancer. They aim to accurately assess performance by collecting significant factors such as inference time, network response time, payload preprocessing and packing time, and investment time. Due to the limited number of available streaming HPC inference benchmarks, iBench is a promising tool for evaluating network I/O. Current limitations of iBench include the inability to represent many ML models and deployable HPC edge devices. Brewer et al. (Brewer et al., 2020a) evaluated iBench, noting several key takeaways, including the impact on performance when using different client batch sizes of , , or , and processing requests asynchronously, the importance of topology-aware GPU-TPU and CPU-GPU communication, and the significance of CPU-GPU in inference compared to training. The evaluation concludes that linear scaling can be achieved if multiple client requests are processed asynchronously and the amount of data sent does not saturate the network.
@@
While benchmarking tools allow ML developers to emulate various ML workload I/O patterns, profiling tools that enable developers to analyze the HPC I/O stack in its entirety are crucial for pinpointing potential bottlenecks. TensorFlow offers the TensorFlowProfiler, released in version , capable of profiling host-side and GPU tracing. However, it only provides information at the TensorFlow level, not capturing lower-level I/O operations such as POSIX or STDIO. Chien et al. (et al, 2020a) created tf-Darshan, a fine-grained I/O profiling tool. It relies on Darshan (et al, 2016c), a highly popular HPC I/O profiler and tracer. Tf-darshan augments Darshan, enabling the runtime extraction and analysis of Darshan’s internal data structures. Runtime extraction is accomplished through an augmentation of the Darshan library. It adds extraction functions that return its internal buffers that store information, such as record counters for I/O operations, to individual files. This gives ML applications the ability to perform runtime analysis, enabling I/O optimization techniques such as specialized online I/O schedulers to perform runtime optimizations. Tf-darshan also offers a visualization tool that allows users to understand the I/O performance of their ML workloads intuitively. Current limitations include exclusive TensorFlow integration and a moderate overhead of - .